# How to Correctly Report LLM-as-a-Judge Evaluations

### v2

@@ Abstract
Large language models (LLMs) are widely used as scalable evaluators of model responses in lieu of human annotators. However, imperfect sensitivity and specificity of LLM judgments induce bias in naive evaluation scores. We propose a simple plug-in framework that corrects this bias and constructs confidence intervals accounting for uncertainty from both the test dataset and a human-evaluated calibration dataset, enabling statistically sound and practical LLM-based evaluation. Building on this framework, we introduce an adaptive calibration strategy for constructing the calibration dataset to reduce uncertainty in the estimated score. Notably, we characterize the regimes in which LLM-based evaluation within our framework produces more reliable estimates than fully human evaluation. Moreover, our framework is more robust to distribution shift between the test and calibration datasets than existing approaches.
Machine Learning, ICML
@@ 1 Introduction
The use of large language models (LLMs) as judges provides a cheap, scalable alternative to human evaluation for various tasks like grading factual accuracy, assessing code quality or detecting harmful content (zheng2023judging; liu2023g; wang-etal-2023-chatgpt; li-etal-2025-generation; gu2024survey). A common practice is to report the proportion of responses that an LLM judges as ‘correct’, denoted by $\hat{p}$. However, directly reporting raw LLM judgment scores is statistically problematic (bross1954misclassification; schwartz1985neglected; forman2005counting; angelopoulos2023prediction; boyeau2025autoeval; fraser2024estimating; albinet2025llmjudge). Because LLM judgments are inherently noisy, $\hat{p}$ generally deviates from the true accuracy and fails to provide a reliable estimate (wang-etal-2024-large-language-models-fair; koo-etal-2024-benchmarking; huang-etal-2025-empirical).
@@
Figure 1: LLM judgments with error rates $1-q_{1}$ and $1-q_{0}$, where $q_{1}$ and $q_{0}$ are LLM’s sensitivity and specificity, respectively.
@@
In extreme cases, such as when $q_{1}=1$ and $q_{0}=0$, the LLM judges every response as ‘correct’. Consequently, the average of the LLM’s binary judgments $\hat{p}$ deviates from the true accuracy $\theta$, becoming equal to $1$ regardless of $\theta$. This illustrates that the raw judgment score can be biased.
@@
implying positive bias at low $\theta$ and negative bias at high $\theta$ (see Section 5 for details). This is illustrated in Figure 2(a) for an LLM with $q_{1}=0.9$ and $q_{0}=0.7$. $\mathbb{E}[\hat{p}$] overestimates $\theta$ when $\theta<0.75$ (blue line) and underestimates it when $\theta>0.75$ (red line), which arises from the two underlying judgment errors (green arrows). A high probability of wrongly rejecting a ‘correct’ response (large $1-q_{1}$) induces negative bias at high accuracies, whereas a high error probability of wrongly accepting an ‘incorrect’ response (large $1-q_{0}$) induces positive bias at low accuracies.
(a) Raw LLM-based judgment score $\hat{p}$
(b) Our adjustment method
(c) Bias-adjusted accuracy $\hat{\theta}$
Figure 2: Bias and its adjustment in LLM-based judgment under imperfect LLM evaluators ($q_{1}=0.9$ and $q_{0}=0.7$). (a) When the true accuracy $\theta$ is low ($\theta<0.75$), the expected value of the raw LLM-based judgment score $\mathbb{E}[\hat{p}$] overestimates $\theta$, whereas when $\theta$ is high ($\theta>0.75$), it underestimates $\theta$. (b) By accounting for the sensitivity $q_{1}$ and specificity $q_{0}$ of the LLM judge, which can be estimated from a calibration dataset with true labels, we obtain the bias-adjusted estimator $\hat{\theta}$ along with its confidence interval (CI). (c) The resulting estimator $\hat{\theta}$ is unbiased when the true values of $q_{0}$ and $q_{1}$ are known or when a calibration dataset of large size is available. A plug-in Python implementation of this procedure is provided in https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting.
@@
This issue is not merely theoretical. With the increasing adoption of LLM-based evaluation, reported improvements may be driven by bias induced by judgment errors rather than true model gains. Since different evaluation procedures can induce biases of different magnitude and direction, some apparent advances in the literature may arise from evaluation artifacts rather than genuine improvements. This motivates careful comparisons and the use of calibrated judges when interpreting past findings, and highlights the need for a principled method for bias adjustment.
@@
Fortunately, this bias can be corrected. When the sensitivity $q_{1}$ and specificity $q_{0}$ are known, a classical result (rogan1978estimating) provides an exact adjustment. Even when they are unknown, they can be estimated from a calibration dataset with human-evaluated labels (buonaccorsi2010measurement), and the resulting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ can be substituted into the correction formula. This adjustment produces the bias-adjusted estimator $\hat{\theta}$ in Figure 2(c) by correcting bias.
Correcting this bias, however, is only part of the problem. LLM-as-a-judge evaluation involves two sources of uncertainty: (i) randomness arising from the test dataset, which affects the judgment score $\hat{p}$, and (ii) randomness from the calibration dataset, which affects $\hat{q}_{0}$ and $\hat{q}_{1}$. A principled confidence interval must incorporate both components; yet prior discussion on LLM-as-a-judge evaluation has focused largely on bias (fraser2024estimating; albinet2025llmjudge), offering no method for constructing valid intervals.
@@
This work provides a statistical framework for LLM-as-a-judge evaluation, as outlined in Figure 2(b). Our key contributions are summarized below:
@@
  • In Section 4, we introduce an estimator for the true accuracy $\theta$ that corrects bias in LLM-based judgments and derive confidence intervals that account for uncertainty from both the test and calibration datasets. We further propose an adaptive strategy for allocating calibration samples that reduces confidence interval length.
  • In Section 6, we characterize the regimes in which LLM-as-a-judge evaluation within our framework yields lower-variance estimates of $\theta$ than direct human evaluation. In Section 7, we show that our framework is more robust to distribution shift between the test and calibration datasets than existing approaches.
@@
  • In Section 5, we provide theoretical justification for our method, and in Section 8, we validate this theory through extensive Monte Carlo simulations and real-world evaluations on the Chatbot Arena benchmark.
@@ 2 Related Work
Recent efforts have sought to formalize the statistical reliability of language model evaluations. While frameworks for calculating error bars and confidence intervals in benchmarks exist (miller2024adding), they assume that LLM evaluators provide ground-truth labels. However, this assumption does not hold in the LLM-as-a-judge setting, where imperfect sensitivity and specificity introduce bias into LLM-based judgment scores. Recent works have examined these biases across various contexts, including natural language inference and test-time compute scaling (godbole-jia-2025-verify; mukherjee2025test; feng2025we). Aligned with this line of work, we propose a bias-adjusted estimator together with statistically sound confidence intervals.
@@
To address bias resulting from imperfect verifiers (e.g., diagnostic or screening tests, or LLM-based judges), rogan1978estimating established an adjustment to correct bias in prevalence estimates given known sensitivity and specificity. This line of work has been extended by more general frameworks such as Prediction-Powered Inference (angelopoulos2023prediction; angelopoulos2023ppi++; zrnic2024cross; broska2025mixed; boyeau2025autoeval), which estimates and corrects bias using a dataset with true labels, as well as calibration-based estimator methods (buonaccorsi2010measurement; kloos2021comparing; meertens2022improving), as detail in Section 7. While each approach offers distinct advantages, our method builds on rogan1978estimating; lang2014confidence, adopting distributional assumptions that are robust to shifts between calibration and test datasets and well-suited to the LLM-as-a-judge setting.
@@ 3 Problem Setup: LLM-as-a-Judge
Evaluation Objective.
We consider the problem of evaluating responses using human judgment. For example, each evaluation instance may consist of a response produced by a given model [^1] The model producing the response may be an LLM, but rule-based or statistical models are also possible., together with the corresponding input, such as a question. We assume that humans can assess whether a given response is ‘correct’ or ‘incorrect’. This assessment is formalized by a ground-truth labeling function $z:\mathcal{X}\to\{0,1\}$, where $z(x)=1$ indicates that humans judge the response in instance $x$ to be ‘correct’, and $z(x)=0$ otherwise. Applying this function to a random instance $X$ induces a binary random variable $Z:=z(X)$.
@@
Let $n]:=\{1,\cdots,n\}$, where $n$ denotes the test set size. Given a test dataset $\{x_{i}\}_{i\in[n]}$ sampled i.i.d. from $\mathbb{P}$}, the LLM produces predictions $\hat{z}_{i}:=f_{\mathrm{LLM}}(x_{i})$. The quantity typically reported in practice is the empirical fraction of instances judged as ‘correct’ by the LLM:
@@
This estimator targets the probability $p:=\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}=1)$ that the LLM judges a randomly drawn test instance as ‘correct’.
Mismatch Between LLM and Human Judgments.
The LLM’s judgment $\hat{Z}$ does not necessarily coincide with the human-evaluated true label $Z$. In particular, the LLM may incorrectly reject responses that are truly ‘correct’ or accept responses that are truly ‘incorrect’. These two types of errors are captured by the following parameters:
$q_{1}:=\mathop{\rm Pr}\nolimits(\hat{Z}=1\!\mid\!Z=1),\,\,\,q_{0}:=\mathop{\rm Pr}\nolimits(\hat{Z}=0\!\mid\!Z=0),\,\,$ (4)
which correspond to the sensitivity (true positive rate) and specificity (true negative rate) of the LLM judge, respectively (forman2008quantifying; lang2014confidence). Unless the LLM is perfectly accurate (i.e., $q_{0}=q_{1}=1$), the naive estimator $\hat{p}$ in Eq. (3) is generally a biased estimator of $\theta$.
@@
Let $\mathbb{Q}$} denote the distribution over calibration instances. Unlike the test dataset, each calibration instance is evaluated by humans, so that the ground-truth variable $Z$ and the corresponding LLM judgment $\hat{Z}$ are observable.
Let $m$ denote the calibration sample size, and let $m_{1}$ and $m_{0}$ denote the numbers of calibration instances with $z_{j}=1$ and $z_{j}=0$, respectively, where $m=m_{0}+m_{1}$. The index $j$ is used to distinguish calibration instances from test instances indexed by $i$. Using this dataset, we estimate the sensitivity $q_{1}$ and specificity $q_{0}$ of the LLM judge in Eq. (4) as
$\hat{q}_{1}$ $:=\frac{1}{m_{1}}\sum_{j\in[m]}\mathbf{1}\{\hat{z}_{j}=1,z_{j}=1\},$ (5)
$\hat{q}_{0}$ $:=\frac{1}{m_{0}}\sum_{j\in[m]}\mathbf{1}\{\hat{z}_{j}=0,z_{j}=0\},$ (6)
where $\mathbf{1}\{\cdot\}$ denotes the indicator function. Throughout this work, we allow the distribution of $Z$ under $\mathbb{P}$} and $\mathbb{Q}$} to differ.
@@
Because the LLM judge is imperfect, the naive estimator $\hat{p}$ generally satisfies $\mathbb{E}[\hat{p}]\neq\theta$, where $\theta$ is the true accuracy defined in (2). Moreover, existing LLM-as-a-judge evaluations typically report only this point estimate, without quantifying uncertainty through confidence intervals. As a result, reported accuracies may appear precise even when they are statistically unreliable.
Our objective is therefore twofold: (i) to construct a bias-adjusted estimator of the true accuracy $\theta$, and (ii) to provide statistically sound confidence intervals that reflect uncertainty arising from both the test and calibration datasets.
@@ 4 Method to Correctly Report LLM-as-a-Judge Evaluations
In this section, we present a bias-adjusted estimator and a confidence interval for LLM-as-a-judge evaluations. We further analyze how sample sizes affect these intervals and propose an adaptive allocation strategy for constructing the calibration dataset that reduces confidence interval length.
Mitigating Bias on Point Estimator.
We begin with the setting in which the sensitivity $q_{1}$ and specificity $q_{0}$ in Eq. (4) are known. In this case, an unbiased estimator of the true accuracy $\theta$ in (2) is given by
$\hat{\theta}\,\big|\,q_{0},q_{1}=\frac{\hat{p}+q_{0}-1}{q_{0}+q_{1}-1},$ (7)
where derivations in this section are deferred to Section 5.
In realistic settings, these accuracies are unknown and must be estimated from a calibration dataset with human-evaluated labels. Substituting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ into (7) gives the bias-adjusted estimator (rogan1978estimating):

### v3

@@ Abstract
Large language models (LLMs) are widely used as scalable evaluators of model responses in lieu of human annotators. However, imperfect sensitivity and specificity of the LLM judges induce bias in naive evaluation scores. We propose a simple plug-in framework that corrects this bias and enables statistically principled uncertainty quantification. Our framework constructs confidence intervals that account for uncertainty from both the test dataset and a human-labeled calibration dataset. Additionally, it uses an adaptive strategy to allocate calibration samples for tighter intervals. Importantly, we characterize parameter regimes defined by the true evaluation score and the LLM judge’s sensitivity and specificity in which our LLM-based evaluation yields more reliable estimates than human-only evaluation. Moreover, we show that our framework remains unbiased under distribution shift between the test and calibration datasets, in contrast to existing approaches.
LLM, evaluation, LLM-as-a-judge, bias, correction, confidence interval, calibration, unbiased estimation, statistical inference
@@ 1 Introduction
The use of large language models (LLMs) as judges provides a cheap, scalable alternative to human evaluation for various tasks such as grading factual accuracy, assessing code quality or detecting harmful content (Zheng et al., 2023; Liu et al., 2023; Wang et al., 2023; Li et al., 2025; Gu et al., 2024). A common practice is to report the proportion of responses that an LLM judges as ‘correct’, denoted by $\hat{p}$. However, directly reporting raw LLM judgment scores is statistically problematic (Bross, 1954; Schwartz, 1985; Forman, 2005; Angelopoulos et al., 2023a; Boyeau et al., 2025; Fraser, 2024; Albinet, 2025). Because LLM judgments are inherently noisy, $\hat{p}$ generally deviates from the true accuracy and fails to provide a reliable estimate (Wang et al., 2024; Koo et al., 2024; Huang et al., 2025).
@@
Figure 1: LLM judgments with error rates $1-q_{1}$ and $1-q_{0}$, where $q_{1}$ and $q_{0}$ are the LLM’s sensitivity and specificity, respectively.
@@
(a) Raw LLM-based estimator $\hat{p}$
(b) Our bias-correction method
(c) Bias-corrected estimator $\hat{\theta}$
Figure 2: Bias and its adjustment in LLM-based judgment under imperfect LLM evaluators ($1-q_{0}=0.3$ and $1-q_{1}=0.1$). (a) When the true accuracy $\theta$ is low ($\theta<0.75$), the expected value of the raw LLM-based estimator $\mathbb{E}[\hat{p}$] overestimates $\theta$, whereas when $\theta$ is high ($\theta>0.75$), it underestimates $\theta$. (b) By accounting for the LLM judge’s sensitivity $q_{1}$ and specificity $q_{0}$, we obtain the bias-corrected estimator $\hat{\theta}$ and its confidence interval (CI). We estimate $q_{1}$ and $q_{0}$ using a calibration dataset with true labels from human evaluators paired with LLM judgments. Here, $n$, $m_{0}$, and $m_{1}$ denote the test-set size, the number of calibration samples whose true label is incorrect (e.g., $4+7=3$), and the number of calibration samples whose true label is correct (e.g., $4+7=11$), respectively. (c) The resulting estimator $\hat{\theta}$ is unbiased when the true values of $q_{0}$ and $q_{1}$ are known, or when a sufficiently large calibration dataset is available. A plug-in Python implementation of this procedure is provided at https://github.com/UW-Madison-Lee-Lab/LLM-judge-reporting.
As an illustration of how the raw LLM judgment scores can be biased, we consider the extreme scenario where $q_{1}=1$ and $q_{0}=0$. In this case, the LLM will judge every response as ‘correct’ and consequently, the average of the LLM’s judgments $\hat{p}$ will be equal to $1$ regardless of the true $\theta$.
@@
implying positive bias at low $\theta$ and negative bias at high $\theta$ (see Section 5 for details). This is illustrated in Figure 2(a) for an LLM with $1-q_{0}=0.3$ and $1-q_{1}=0.1$. Specifically, $\mathbb{E}[\hat{p}$] overestimates $\theta$ when $\theta<0.75$ (blue line) and underestimates it when $\theta>0.75$ (red line), which arises from the two underlying judgment errors (green arrows). A high probability of incorrectly rejecting a ‘correct’ response (large $1-q_{1}$) induces negative bias at high accuracies, whereas a high probability of incorrectly accepting an ‘incorrect’ response (large $1-q_{0}$) induces positive bias at low accuracies.
@@
This issue is not merely theoretical. With the increasing adoption of LLM-based evaluation, reported improvements may be driven by bias induced by judgment errors rather than true model gains. Since different evaluation procedures can induce biases of different magnitude and direction, some apparent advances in the literature may arise from evaluation artifacts rather than genuine improvements. This highlights the importance of careful comparisons and the use of calibrated judges when interpreting past findings, and motivates the need for a principled method for bias adjustment.
@@
Fortunately, this bias can be corrected. When the sensitivity $q_{1}$ and specificity $q_{0}$ are known, a classical result (Rogan and Gladen, 1978) provides an exact adjustment. When they are unknown, they can be estimated from a calibration dataset with human-evaluated labels, and the resulting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ can be substituted into the correction formula. This leads to the bias-corrected estimator $\hat{\theta}$ in Figure 2(c).
More broadly, a variety of bias-correction methods have been studied for imperfect evaluators. In particular, prediction-powered inference (Angelopoulos et al., 2023a, b) provides a general framework that extends beyond categorical outcomes and can yield estimators of the true accuracy $\theta$ with lower variance; see Kloos et al. (2021); Chen et al. (2026). However, we show that existing methods can become biased under distribution shift between the test and calibration datasets, which can arise in LLM-as-a-judge settings (Jung et al., 2024). In contrast, the estimator $\hat{\theta}$ based on the method of Rogan and Gladen (1978) remains unbiased under such shifts. Accordingly, we adopt this estimator and focus our analysis on $\hat{\theta}$. Moreover, since the estimator $\hat{\theta}$ may exhibit higher variance than existing alternatives, we propose an adaptive strategy for allocating calibration samples that reduces its variance and, consequently, shortens the confidence interval for $\theta$ under our method.
@@
In this paper, we provide a statistical framework for correctly reporting LLM-as-a-judge evaluations, as outlined in Figure 2(b). We summarize our key contributions below:
@@
  • In Section 4, we introduce the bias-corrected estimator for the true accuracy $\theta$, derive the confidence interval, and further propose an adaptive strategy for allocating calibration samples that reduces its length. In Section 5, we establish theoretical guarantees for our method.
  • In Section 6, we characterize parameter regimes defined by $q_{1}$, $q_{0}$, and $\theta$ in which our estimator has smaller variance than that obtained from human-only evaluation.
@@
  • In Section 7, we validate our method through extensive Monte Carlo simulations and real-world LLM-as-a-judge evaluations on the Chatbot Arena benchmark.
  • In Section 8, we show that estimators from existing bias-correction methods can still be biased under distribution shift between the test and calibration datasets, whereas our estimator remains unbiased under such shifts.
@@ 2 Related Work
Recent efforts have sought to formalize the statistical reliability of language model evaluations. While frameworks for calculating error bars and confidence intervals in benchmarks exist (Miller, 2024), they assume that LLM evaluators provide true labels. However, this assumption does not hold in the LLM-as-a-judge setting, where imperfect judges introduce bias into LLM-based scores. Recent works have examined these biases across various contexts, including natural language inference and test-time compute scaling (Godbole and Jia, 2025; Mukherjee et al., 2025; Feng et al., 2025). Aligned with this line of work, we propose a bias-corrected estimator together with statistically sound confidence intervals.
@@
To address bias arising from imperfect evaluators (e.g., diagnostic or screening tests, or LLM-based judges), Rogan and Gladen (1978) proposed an adjustment that corrects prevalence estimates when sensitivity and specificity are known. Subsequent work has developed more general frameworks, most notably Prediction-Powered Inference (Angelopoulos et al., 2023a, b; Zrnic and Candès, 2024; Broska et al., 2025; Boyeau et al., 2025), which leverages a small set of true labels to improve estimation, as well as conditional calibration estimators (Buonaccorsi, 2010; Kloos et al., 2021; Meertens et al., 2022); see Section 8 for details. While these methods offer distinct advantages, such as variance reduction (Kloos et al., 2021; Chen et al., 2026), our approach builds on Rogan and Gladen (1978); Lang and Reiczigel (2014), adopting distributional assumptions that remain valid under shifts between the calibration and test datasets and are particularly well-suited to the LLM-as-a-judge setting.
@@ 3 Problem Setup: LLM-as-a-Judge
We consider the problem of evaluating responses based on human judgment. For example, each instance consists of a question and a corresponding response produced by a given model [^1] The model producing the response may be an LLM, but rule-based or statistical models are also possible.. We assume that humans can assess whether a response is ‘correct’ or ‘incorrect’. This assessment is formalized by a ground-truth labeling function $z:\mathcal{X}\to\{0,1\}$, where $z(x)=1$ indicates that humans judge the response in instance $x$ to be ‘correct’, and $z(x)=0$ otherwise. Applying the function to a random instance $X$ induces a binary random variable $Z=z(X)$.
@@
Let $n]:=\{1,\cdots,n\}$, where $n$ denotes the test-set size. Given a test dataset $\{x_{i}\}_{i\in[n]}$ sampled i.i.d. from $\mathbb{P}$}, the LLM produces predictions $\hat{z}_{i}:=f_{\mathrm{LLM}}(x_{i})$. The quantity reported in practice is the empirical fraction of instances judged as ‘correct’ by the LLM:
@@
This estimator targets the probability $p:=\mathop{\rm Pr}\nolimits(\hat{Z}=1)$ that the LLM judges a test instance as ‘correct’.
However, the LLM’s judgment $\hat{Z}$ does not necessarily coincide with the human-evaluated true label $Z$. We characterize the LLM judge using the following parameters:
$q_{1}:=\mathop{\rm Pr}\nolimits(\hat{Z}=1\!\mid\!Z=1),\;q_{0}:=\mathop{\rm Pr}\nolimits(\hat{Z}=0\!\mid\!Z=0).\;\;$ (4)
which correspond to the sensitivity (true positive rate) and specificity (true negative rate) of the LLM judge, respectively (Forman, 2008; Lang and Reiczigel, 2014). Consequently, the LLM may incorrectly reject responses that are truly ‘correct’ with probability $1-q_{1}$, or accept responses that are truly ‘incorrect’ with probability $1-q_{0}$. Unless the LLM is perfectly accurate (i.e., $q_{0}=q_{1}=1$), the naive estimator $\hat{p}$ in (3) is generally a biased estimator of $\theta$.
@@
Let $\mathbb{Q}$} denote the distribution over calibration instances. Unlike the test dataset, each calibration instance is evaluated by humans, so that both the true label $Z$ and the corresponding LLM judgment $\hat{Z}$ are observable.
Let $m$ denote the calibration-set size, and let $m_{1}$ and $m_{0}$ denote the numbers of calibration instances with $z_{j}=1$ and $z_{j}=0$, respectively, so that $m=m_{0}+m_{1}$. The index $j$ distinguishes calibration instances from test instances indexed by $i$. Using the calibration dataset, we estimate the sensitivity $q_{1}$ and specificity $q_{0}$ of the LLM judge in (4) as
$\hat{q}_{1}\!:=\!\!\sum_{j\in[m]}\!\frac{\mathbf{1}\{\hat{z}_{j}=1,z_{j}=1\}}{m_{1}},\;\hat{q}_{0}\!:=\!\!\sum_{j\in[m]}\!\frac{\mathbf{1}\{\hat{z}_{j}=0,z_{j}=0\}}{m_{0}},\!\!\!$
where $\mathbf{1}\{\cdot\}$ denotes the indicator function. In the main analysis, we assume $\mathbb{P}$$\mathbb{Q}$}. In Section 8, we relax this assumption and consider distribution shift, where $\mathbb{P}$$\neq{\mathbb{Q}}$.
@@
Because the LLM judge is imperfect, the naive estimator $\hat{p}$ is generally biased for the true accuracy $\theta$ defined in (2), i.e., $\mathbb{E}[\hat{p}]\neq\theta$. Moreover, existing LLM-as-a-judge evaluations typically report this point estimate without quantifying uncertainty via confidence intervals.
Our objective is therefore twofold: (i) to construct a bias-corrected estimator of the true accuracy $\theta$, and (ii) to provide statistically sound confidence intervals that reflect uncertainty arising from both the test and calibration datasets.
@@ 4 Method to Correctly Report LLM-as-a-Judge Evaluations
In this section, we present a bias-corrected estimator $\hat{\theta}$ and a confidence interval for LLM-as-a-judge evaluations. We further propose an adaptive allocation strategy for constructing the calibration dataset that reduces the confidence interval length. Derivations and theoretical guarantees for the proposed method are provided in Section 5.
4.1 Mitigating Bias in Point Estimator
We begin with the setting in which the sensitivity $q_{1}$ and specificity $q_{0}$ in (4) are known. In this case, an unbiased estimator of the true accuracy $\theta$ in (2) is given by
$\hat{\theta}\,\big|\,q_{0},q_{1}=\frac{\hat{p}+q_{0}-1}{q_{0}+q_{1}-1}.$ (6)
In realistic settings, these parameters $q_{1}$ and $q_{0}$ are unknown and must be estimated from a calibration dataset. Substituting estimates $\hat{q}_{1}$ and $\hat{q}_{0}$ into (6) gives the bias-corrected estimator (Rogan and Gladen, 1978; Lang and Reiczigel, 2014):